OpenStudy (anonymous):

Simplify. Assume all variables represent positive real numbers. ^3sqrt81x^8

OpenStudy (anonymous):

27x^6

OpenStudy (campbell_st):

\[\sqrt[3]{81x^8} = \sqrt[3]{27 \times3 \times x^6 \times x^2} = 3x^2\sqrt[3]{3x^2}\]

OpenStudy (campbell_st):

27 = 3^3 x^6 = (x^2)^3